OpenStudy (anonymous):

How do I know this is true? 2x -4 = 2x - 4

OpenStudy (anonymous):

its true because its the commutative property i believe

OpenStudy (anonymous):

You can use login to see that it's the same but here is the answer: 2x−4=2x−4 So 2x−2x= -4 + 4 0=0 So it's true.

OpenStudy (anonymous):

thanks for the medal

OpenStudy (anonymous):

I meant logic instead of login. Sorry.

OpenStudy (anonymous):

Thanks guys :)

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

no problem.